zarkam21:

𝑓(𝑥) = 𝑥^2 + 1, 𝑔(𝑥) = √𝑥. Determine the derivative of 𝑖) 𝑓(𝑔(𝑥)), 𝑖𝑖) 𝑔(𝑓(𝑥)), 𝑖𝑖𝑖) 𝑓(𝑓(𝑥)), and 𝑖𝑣) 𝑔(𝑔(𝑥)) with respect to 𝑥

zarkam21:

@Vocaloid

zarkam21:

Just need a better explanation

Vocaloid:

well i) is just asking for the derivative of 𝑓(𝑔(𝑥)), right? so you just take g(x) = sqrt(x) and plug that into the x-value for f(x), then take the derivative of the result

zarkam21:

would the exponent 2 go in the sqrt too or no

mhchen:

No it'd go outside the square root. f(x) = x^2 g(x) =√x f(g(x)) = f(√x) = (√x)^2 + 1 get it?

zarkam21:

okay is that it

mhchen:

No you gotta take the derivative of that. You know how to take the derivatives right?

zarkam21:

um kind of can you just clear up the concept

mhchen:

Yeah so the first question asks you to take the derivative of f(g(x)) and you know f(g(x)) is (√x)^2+1 which can be simplified to x + 1 When you take the derivative of something, you basically use this formula: \[\frac{d}{dx}ax^b =(b*a)x^{b-1}\] so x + 1 can be rewritten as \[1x^{1}+1x^0\] can you find the derivative of that?

zarkam21:

x^-1

mhchen:

oh shoot wait I made a mistake. Sorry. 1x^1 + 1 Since 1 is a constant, the derivative of constants is always 0. so the derivative is actually \[(1*1)x^{1-1}+0\] which is just 1.

mhchen:

I can give you some more examples of derivatives: derivative of \[\frac{d}{dx}(5x^3+2x^2+x+4) = (3*5)x^2+(2*2)x+(1) = 15x^2+4x+1\]

Vocaloid:

idk if it'll make this more/less confusing but generally if you have something like A^B you bring the exponent down as a coefficient to get B * A^B then reduce the exponent by 1 to get B* A^(B-1) then multiply by the derivative of the base (A) --> B * A^(B-1) * derivative of A

Vocaloid:

it just takes a bit of time to learn the pattern that's all

zarkam21:

I'm taking notes on this so we can continue in like 5 mins

Vocaloid:

ok

zarkam21:

okay so derivative of sqrtx^2+1 is 2x^2?

mhchen:

Well actually sqrtx^2 + 1 = x + 1 right?

mhchen:

so derivative of x + 1 is ?

mhchen:

It's just 1. Derivative of x + 1 is just 1. Derivative of x is 1. Derivative of 1 (or any constant) is 0. So it becomes 1 + 0. Do you understand why?

zarkam21:

Yes I do understand why

zarkam21:

Could you just explain what happens to the exponent though

mhchen:

Yeah so the exponent gets brought down in front of x as a multiplier, and the exponent on top decreases by 1. x^3 = 3x^2 x^2 = 2x 4x^2 = (4*2)x = 8x

zarkam21:

No I mean in the sqrtx^2+1 problem

Vocaloid:

taking the square root of something will undo a square

Vocaloid:

sqrt(A^2) = A

zarkam21:

Got it

mhchen:

oh yeah. you can also think of it like this: \[\sqrt{x} = x^{\frac{1}{2}}\] and \[(x^a)^{b} = x^{a*b}\] so \[\sqrt{x}^2 = (x^{\frac{1}{2}})^2 = x^{\frac{1}{2}*2} = x\]

Vocaloid:

anyway that's the first derivative done; derivative of 𝑖) 𝑓(𝑔(𝑥)) = 1 according to our calculations if you want to start the second one 𝑖𝑖) 𝑔(𝑓(𝑥)) I'm ready when you are

zarkam21:

OKay so f(x)=x^2+1 g(x)=sqrtx g(f(x)) g(sqrt^2)

zarkam21:

Wait another question :S where did the +1 come from in the first one :/

mhchen:

it came from f(x) = x^2 + 1

zarkam21:

oh right

zarkam21:

OKay so f(x)=x^2+1 g(x)=sqrtx g(f(x)) g(sqrt^2)+1 =1

mhchen:

Well actually. g(f(x)) is g( sqrt^2+1 )

zarkam21:

I thought we were on ii

Vocaloid:

but yeah you need to pay attention to parentheses

zarkam21:

so ii is g( sqrt^2+1 )

Vocaloid:

yes

Vocaloid:

now applying g(x) = sqrt(x) g(f(x)) = g( sqrt^2+1 ) = sqrt(x^2+1) then take the derivative of this

zarkam21:

x?

Vocaloid:

re-write sqrt(x^2+1) as (x^2 + 1)^(1/2) and apply the derivative rule

zarkam21:

1/2(x^2+1)^-1/2

Vocaloid:

almost, you need to also multiply by the derivative of x^2 + 1

zarkam21:

2x+1?

zarkam21:

2x?

zarkam21:

:S

Vocaloid:

x^2 + 1 ---> derivative is 2x (since 1 is just a constant) so (x^2 + 1)^(1/2) derivative ---> (1/2)(2x)(x^2+1)^(-1/2) ---> x * (x^2+1)^(-1/2) = x/sqrt(x^2+1) = your derivative

mhchen:

zarkam have you learned the chain rule before?

zarkam21:

\[\frac{ x }{ \sqrt{x^2+1} }\]

zarkam21:

No , I haven't my professor is reallly not helping any of us which is why me along with other students are struggling

zarkam21:

he goes over problems and says to solve them you use the chain rule. We ask him what the rule is and he says we haven't gotten there yet

zarkam21:

okay for iii it would be f(f(x) f(x^2+1)

zarkam21:

I think :S

mhchen:

Yeah you're right so far. me/voca will help you learn the chain rule

zarkam21:

2x as the derivative

Vocaloid:

f(x^2+1) since f(x) = x^2 + 1, then you need to plug "x^2+1" into f(x) to get (x^2+1)^2 + 1 then take the derivative of that

zarkam21:

2x^2?

zarkam21:

orrr 2(2x)

mhchen:

I think Vocaloid wants you to simplify that before you take the derivative. Just simplify (x^2+1)^2 + 1 first

zarkam21:

(x^2)^2

mhchen:

Well actually \[(x^2+1)^2 + 1 = (x^4 + 2x^2 + 1) + 1 = x^4 + 2x^2 + 2\] right?

zarkam21:

Yes

mhchen:

Alright so find the derivative of \[x^4 + 2x^2 + 2\]

zarkam21:

(4)x^3+(2x2)x+2=4x^3+2x^2

zarkam21:

(4)x^3+(2x2)x+2=4x^3+6x^2

mhchen:

uh, check that last one

mhchen:

4x^3 + (2x2)x = 4x^3 + 4x did you see what you did wrong?

zarkam21:

Yess

mhchen:

okay well there's your answer

mhchen:

See how long it takes to do this? I wanna show you a trick: If you have a function like f(x) = x^2 and you have another function like g(x) = x^3 and you want to find the derivative of f(g(x)) you can rewrite it as f'(g(x)) * g'(x) so f'(x) = 2x and f'(g(x)) = 2(g(x)) = 2(x^3) g'(x) = 3x^2 f'(g(x))g'(x) = (2x^3)(3x^2) does that make sense?

mhchen:

f'(x) means "the derivative of f(x)" btw

zarkam21:

Yeah it makes more sense than what we went over before